The value of a two-digit number is twice as large as the sum of its digits. If the digits were reversed, the resulting number would be 9 less than 5 times the original number. Find the original number.
Answer by stanbon(75887)

Let the number be ab
It's "value" is 10a+b
EQUATION#1:
10a+b=2(a+b)
8a=b
If the digits are reversed the number is ba
Then it's value is 10b+a
EQUATION#2:
10b+a+9=5(10a+b)
5b+9=49a
Make a substitutin for "b", using b=8a.
the 5(8a)+9=49a
a=1
Substitute into b=8a to get b=8
The original number is ab=18
Cheers,
stan H.